Amortize Discount by Interest Method On the first day of its fiscal year, Ebert Company issued $12,000,000 of 5-year, 11% bon3. Second semiannual interest payment, including amortization of discount. Round to the nearest dollar. If an amount box doesOn the first day of its fiscal year, Ebert Company issued $12,000,000 of 5-year, 11% bonds to finance its operations. Interest is payable semiannually. The bonds were issued at a market (effective) interest rate of 12%, resulting in Ebert receiving cash of $11,558,459. The company uses the interest method.
Answer #1
a
1 Cash 11558459
Discount on Bonds payable 441541
       Bonds payable 12000000
2 Interest expense 693508 =11558459*12%/2
      Discount on Bonds payable 33508
      Cash 660000 =12000000*11%/2
3 Interest expense 695518 =(11558459+33508)*12%/2
      Discount on Bonds payable 35518
      Cash 660000 =12000000*11%/2
b
Annual interest paid 1320000 =660000*2
Discount amortized 69026 =33508+35518
Interest expense for first year 1389026